# Maintenance Optimization for Asset Networks with Unknown Degradation Parameters

### v2

@@ Abstract
We consider the key practical challenge of multi-asset maintenance optimization in settings where degradation parameters are heterogeneous and unknown, and must be inferred from degradation data. To address this, we propose scalable methods suitable for complex asset networks. Degradation is modeled as a stochastic shock process, and real-time data are continuously incorporated into estimation of shock rates and magnitudes via a Bayesian framework. This constitutes a partially observable Markov decision process formulation, from which we analytically derive monotonic policy structures. Moreover, we propose an open-loop feedback approach that enables policies trained via deep reinforcement learning (DRL) in a simulation environment with access to the true parameters to remain effective when deployed with real-time Bayesian point estimates instead. Complementing this, we develop a Bayesian Markov decision process (BMDP) framework wherein the agent maintains and updates posterior distributions during deployment. This formulation captures the evolution of parameter uncertainty over time, thereby facilitating the training of scalable DRL-based policies that adapt as additional data become available.
We validate our approach through experiments on synthetic asset networks and a real-world case involving interventional X-ray system filaments. We find that the proposed DRL methods consistently outperform traditional heuristics across various scenarios. The policies trained for the BMDP perform well even when priors must be estimated from historical data, and remain effective in networks with high asset heterogeneity. Knowledge of true degradation parameters yields only marginal cost benefits, underscoring the ability of our approach to make effective decisions under limited information on degradation processes.
@@
Maintenance , Optimization , Bayesian inference , Deep reinforcement learning
@@ 1 Introduction
Optimizing the maintenance of an installed base of capital goods, such as wind turbine parks, medical imaging equipment, and wafer steppers, is critical for ensuring continuous availability and operational efficiency. The unavailability of such assets imposes significant costs, with an estimated annual economic impact of around \$50\\$50 billion (Coleman et al., 2017; Wall Street Journal Custom Studios, 2017). Nearly half of this unavailability results from component failures, which not only cause substantial financial losses but also pose safety hazards and disrupt essential supply chains. Minimizing these failures—particularly within sectors such as energy, healthcare, and semiconductors—is therefore essential to reduce downtime costs and mitigate safety risks.
@@
To achieve maximum availability and minimize unnecessary maintenance costs, condition-based maintenance (CBM) has become a prominent proactive maintenance strategy. CBM uses sensors to continuously monitor the degradation of components, triggering maintenance actions only when failure becomes imminent. While CBM may significantly reduce unnecessary interventions and costly downtime, effectively utilizing real-time degradation data to optimally schedule maintenance across an asset network remains a non-trivial challenge.
@@
The challenge of optimizing CBM in real-world settings arises from two interrelated issues: asset dependencies and component heterogeneity. Asset dependencies, such as economic dependencies arising from joint maintenance setup costs, shared spare parts and limited maintenance resources, significantly influence maintenance scheduling across networks of interconnected assets (Olde Keizer et al., 2017). Despite their practical significance, such dependencies are often ignored in traditional CBM optimization approaches due to the analytical complexity that results from simultaneously tracking the condition of multiple assets.
@@
Component heterogeneity presents an additional challenge, as even nominally identical components can exhibit significant variability in degradation behavior due to differences in manufacturing, environmental conditions, or operational usage. Traditional CBM typically assumes known degradation parameters and does not incorporate real-time parameter learning from observed degradation data (De Jonge and Scarf, 2020; Arts et al., 2025). Increasingly, scholars address the challenge of real-time learning by integrating Bayesian inference into maintenance optimization to dynamically update beliefs about component degradation parameters. While promising, such Bayesian models have primarily been studied in single-asset contexts (Elwany et al., 2011; Kim and Makis, 2013; Chen et al., 2015; Van Oosterom et al., 2017; Drent et al., 2023), where they remain analytically tractable. Consequently, the case involving multiple interdependent assets, despite being of considerable practical importance, remains underexplored.
@@
We contribute a CBM model for asset networks with unknown degradation parameters that incorporates economic dependencies arising from shared setup costs—such as the fixed expenses incurred when deploying maintenance crews to service an offshore wind farm. Component degradation is modeled using stochastic shock models (Esary and Marshall, 1973; Sobczyk, 1987). We assign prior distributions to represent uncertainty about key degradation parameters—specifically, the shock occurrence rate and the distributional parameter for damage incurred per shock. Conjugate priors facilitate analytically tractable Bayesian updating as degradation data accumulate, following the single-asset framework in Drent et al. (2023). We derive analytical monotonicity results for optimal maintenance policies in our multi-asset model, with respect to both the degradation levels and the degradation parameters.
@@
Moreover, we contribute methodological innovations and insights that enable the training of policies via deep reinforcement learning (DRL) to optimize CBM for networks of interdependent assets that degrade according to a shock process with unknown parameters, which are progressively inferred from real-time degradation signals. Specifically, we develop two distinct Bayesian simulation environments for a multi-asset setting, each supporting a different DRL approach. The first modeling approach explicitly simulates the true degradation parameters for each component, drawn from a population distribution at the time of component replacement. Subsequent degradation then evolves according to a stochastic shock process governed by these sampled parameters. Since the true degradation parameters are unobservable in practice, this setup constitutes a partially observable Markov decision process (POMDP). To enable deployment of the trained policies, we propose an open-loop feedback approach. Concretely, we collapse the posterior distributions into point estimates, which serve as input to the neural network in place of the true degradation parameters. To complement this first approach, we propose a second modeling approach that assumes the availability of a conjugate prior and, in that case, is equivalent in objective to the first approach—although the two differ in methodology. Instead of simulating the true degradation parameters, this approach explicitly maintains and updates uncertainty by tracking only the posterior distributions of the parameters. These posteriors are continuously updated through real-time degradation observations. This approach constitutes a Bayesian Markov decision process (BMDP), which forms the foundation for the development of DRL-based policies that operate directly on the posterior distributional information available to the asset manager. This marks a significant advancement beyond single-asset models by jointly addressing learning and maintenance optimization in heterogeneous asset networks.
@@
As the training algorithm for our DRL-based policies, we propose the application of specific forms of approximate policy iteration (API) (Temizöz et al., 2025) and compare the resulting policies to relevant benchmarks. In particular, two heuristic policies are adapted as benchmarks and serve as initializations for API: (i) a two-threshold control-limit heuristic that groups maintenance actions, and (ii) an integrated Bayes heuristic known to yield near-optimal performance in the single-asset setting.
@@
Through extensive numerical experiments and a practical case study utilizing real-world degradation data from interventional X-ray (IXR) filaments, we demonstrate the effectiveness and practical applicability of our DRL approaches. To better understand how policy performance varies with the availability of degradation information, we extend the concept of information levels (Da Costa et al., 2023) to our Bayesian setting: Information level $\textbf{L}_{2}$ corresponds to full knowledge of the degradation parameters for each component. In information level $\textbf{L}_{1}$, this parameter information is unknown, but the distribution of degradation parameters is available. Finally, information level $\textbf{L}_{0}$—used only in the case study—represents a setting where even the distributional information must be estimated from historical data.
@@
The remainder of the paper is organized as follows. Section 2 presents an overview of the relevant literature. We provide a detailed model formulation in Section 3. In Section 4, we detail the heuristic solutions and provide a brief overview of the DRL algorithm. We demonstrate the effectiveness of the proposed solutions through a concise simulation study in Section 5. In Section 6, we present the results of the case study on degrading IXR filaments. We offer concluding remarks in Section 7.
@@ 2 Literature review
Our work focuses on the optimization of CBM for asset networks with economic dependencies, unknown degradation parameters inferred from real-time data, and the application of DRL. We next review these three research streams.
@@
Maintenance models may be classified as single-asset or multi-asset (Olde Keizer et al., 2017; De Jonge and Scarf, 2020). Multi-asset models extend single-asset models by incorporating joint maintenance policies, accounting for various dependencies such as economic, structural, stochastic, or resource dependencies. Economic dependence refers to the cost relationship between maintaining multiple assets together versus individually. If combined maintenance is more expensive, it indicates negative economic dependence; if it is less expensive, it indicates positive economic dependence.
@@
Various models for CBM in systems with economic dependence have been developed in recent literature due to their relevance in industrial settings. Wijnmalen and Hontelez (1997) propose a heuristic algorithm that minimizes long-run average costs in a multi-component system by determining simple, component-wise control limits and leveraging coordinated repairs to reduce shared set-up costs. Bouvard et al. (2011) propose a maintenance optimization method for multi-component systems, demonstrated on commercial heavy vehicles, which dynamically groups tasks using updated degradation data. Tian et al. (2011) proposes a CBM strategy for wind farms that leverages neural networks to predict the degradation level and corresponding remaining useful lifetime (RUL) from real-time data. Tian and Liao (2011) develop a CBM policy for multi-component systems using a proportional hazards model and a numerical algorithm to evaluate the cost of the proposed policy.
@@
Zhu et al. (2015) introduce a novel CBM policy for multi-component systems experiencing continuous stochastic degradation, utilizing a joint maintenance interval to reduce shared setup costs. Olde Keizer et al. (2016) develop a dynamic programming model to determine the optimal CBM strategy for systems with both economic dependencies and redundancy, demonstrating that it significantly outperforms heuristic policies and offers insights into the optimal policy structure. Olde Keizer et al. (2018) investigate the trade-off between prompt replacement of failed components and maintenance clustering opportunities in redundant systems with economic dependencies, revealing that heuristic threshold policies may be suboptimal and that misinterpreting load-sharing effects can lead to higher maintenance expenses. Do et al. (2019) present a model for a CBM policy in a two-component system with stochastic dependencies, where the degradation of each component is influenced by the other’s state, and economic dependencies, where combined maintenance activities are shown to be more cost-effective. Similarly, Oakley et al. (2022) propose a CBM policy for multi-component systems that balances the trade-offs between the urgency of replacing failed components due to increased load and the economic benefits of clustering replacements to minimize downtime and maintenance costs. We refer the reader to Zhu et al. (2015, Table 1) for a more extensive summary of CBM models proposed for multi-component systems.
@@
The geographical layout of asset networks often introduces a complex dependency. Abdul-Malak and Kharoufeh (2018) investigate the challenge of optimally replacing multiple stochastically degrading systems within a shared environment using CBM. Soltani et al. (2024) study the problem of optimally maintaining an offshore wind turbine farm under both economic and stochastic dependence due to shared maintenance setup costs and their common environment. They establish monotonicity of the cost function jointly in the degradation level and environmental state, characterize the structure of the optimal replacement policy, and show that sharing maintenance resources is cost-effective. Leppinen et al. (2025) introduce directed graphs to represent the economic and structural dependencies of a multi-component system, including scenarios where maintenance on one component may require the disassembly or maintenance of others, and solve the resulting MDP using a modified policy iteration algorithm to determine the most cost-efficient policy.
@@
Most existing studies on uncertainty in degradation processes typically assume a specific parametric form, with uncertainty modeled through its parameters and captured via a prior distribution. Bayesian inference is then used to update this uncertainty over time—but so far, these approaches have been limited to single-asset systems only. Van Oosterom et al. (2017) develop a POMDP model to incorporate population heterogeneity for maintenance scheduling of a single-asset system that stochastically degrades, however, the population of spare parts consists of multiple indistinguishable types that degrade at varying rates. Elwany et al. (2011) and Si et al. (2018) both study a Wiener degradation process with an unknown drift parameter. In both studies, periodic inspections are used to estimate the parameter, with Bayesian inference applied in the former and maximum likelihood estimation in the latter. These estimation methods, combined with observed degradation levels, guide the decision-making process for maintenance interventions. Flage et al. (2012) consider a single asset subject to a degradation process with unknown parameters. Sequential inspections assess the degradation level, with the timing of the next inspection based on the current degradation level. Uncertainty in the parameters is modeled using a prior distribution and updated in a Bayesian manner. Drent et al. (2023) investigate a single-asset model with one critical component that degrades according to a compound Poisson process with unknown parameters, leveraging real-time degradation data to infer the component’s degradation behavior and adjust decision-making accordingly. They demonstrate that, accounting for component heterogeneity, the optimal policy depends on both the asset’s age and the observed degradation signal, and that integrating learning with decision-making yields significant cost reductions.
@@
DRL-based approaches to maintenance optimization have demonstrated promising results across various settings. Kuhnle et al. (2019) train opportunistic DRL-based maintenance strategies using proximal policy optimization for parallel assets, achieving reductions in downtime and costs compared to traditional strategies such as reactive and time-based maintenance. Zhang and Si (2020) propose a DRL algorithm based on deep Q-network (DQN) to train policies for CBM in multi-component systems with stochastic and economic dependencies, and show in a numerical study that the trained policies outperform benchmarks set by heuristic policies. Mohammadi and He (2022) develop a DRL-based method using double DQN (DDQN) for the joint optimization of maintenance and renewal planning under practical constraints. Using historical inspection and maintenance data to simulate a rail infrastructure environment, they show that the approach generates cost-effective policies that enhance network reliability and safety. Hung et al. (2024) employ DDQN to train DRL-based maintenance policies in a stochastic factory setting characterized by various degradation levels, uncertain repair times and fluctuating machine workloads.
@@
Lee and Mitici (2023) develop an integrated predictive maintenance framework that combines RUL estimation via convolutional neural networks with a DRL-based maintenance policy trained using a soft actor–critic algorithm. The policy triggers maintenance actions based on the estimated RUL distribution and yields significant cost savings and downtime reduction for aircraft turbofan engines. Da Costa et al. (2023) propose a DRL approach based on distributional DQN to minimize maintenance and downtime costs in asset networks serviced by a single maintenance engineer, where degradation levels are only partially observable. Building upon this work, Verleijsdonk et al. (2024) develop DRL-based solutions trained via a form of API—shown to scale better than the aforementioned training algorithms—for industrial asset networks maintained by multiple engineers; however, their approach assumes full observability of all degradation levels.
@@
Recent studies have made important progress (see, e.g., Andriotis and Papakonstantinou (2021)) by integrating Bayesian inference with DRL in the context of maintenance optimization. These efforts focus on inferring the degradation state of components rather than the parameters of the underlying degradation process. Thus, while such studies represent key steps toward maintenance optimization under uncertainty, they typically assume that the parameters of the degradation processes are known and can be used to define the simulation environment in which the DRL-based policy is trained.

### v3

@@ Abstract
We consider the key challenge of maintenance optimization for asset networks where degradation parameters are heterogeneous and unknown, and must be inferred from real-time degradation data. Combining stochastic modeling and Bayesian statistics, we formulate a partially observable Markov decision process that incorporates estimating shock rates and sizes. This formulation allows us to analytically establish that optimal replacement policies are threshold-type in both degradation levels and parameters. Moreover, we propose an open-loop feedback approach that allows policies trained via deep reinforcement learning (DRL) with access to true parameters to remain effective when deployed using real-time Bayesian point estimates. Complementarily, we develop a Bayesian Markov decision process (BMDP) framework in which the agent maintains and updates posterior distributions during deployment, capturing parameter uncertainty evolution and enabling scalable DRL policies that adapt as new data become available. We evaluate our approaches on synthetic data and a real-world case involving interventional X-ray filaments. The proposed DRL methods consistently outperform traditional heuristics. Policies trained for the BMDP remain robust when priors are estimated from historical data and perform well in highly heterogeneous networks. Access to true parameters provides only marginal cost improvements, highlighting the ability of the approach to make effective decisions under limited information.
@@
Condition-based maintenance , Optimization , Bayesian inference , Deep reinforcement learning
@@ 1 Introduction
Optimizing the maintenance of an installed base of capital goods, such as wind turbine parks, medical imaging equipment, and wafer steppers, is critical for ensuring continuous availability and operational efficiency. The unavailability of such assets imposes significant costs, with an estimated annual economic impact of around \$50 billion [Coleman et al., 2017, Wall Street Journal Custom Studios, 2017]. Nearly half of this unavailability results from component failures, which not only cause substantial financial losses but also pose safety hazards and disrupt essential supply chains. Minimizing these failures—particularly within sectors such as energy, healthcare, and semiconductors—is therefore essential to reduce downtime costs and mitigate safety risks.
@@
To achieve maximum availability and minimize unnecessary maintenance costs, condition-based maintenance (CBM) has become a prominent proactive maintenance strategy. In CBM, sensors continuously collect operational data and transmit them in real time. The degradation signal is constructed from these measurements to represent the health of the asset over time, enabling maintenance actions to be triggered only when failure becomes imminent. While CBM may significantly reduce unnecessary interventions and costly downtime, effectively utilizing real-time degradation data to optimally schedule maintenance across an asset network remains a non-trivial challenge.
@@
The challenge of optimizing CBM for real-world asset networks arises from two interrelated issues: asset dependencies and component heterogeneity. Asset dependencies, such as economic dependencies arising from joint maintenance setup costs, shared spare parts and limited maintenance resources, significantly influence maintenance scheduling across networks of interconnected assets [Olde Keizer et al., 2017]. Despite their practical significance, such dependencies are often ignored in traditional CBM optimization approaches due to the analytical complexity that results from simultaneously tracking the condition of multiple assets.
@@
Component heterogeneity presents an additional challenge, as even nominally identical components can exhibit a significant variability in degradation behavior due to differences in manufacturing, environmental conditions, or operational usage. Traditional CBM approaches typically assume known degradation parameters and do not incorporate real-time parameter learning from observed degradation data [De Jonge and Scarf, 2020, Arts et al., 2025]. Increasingly, scholars address the challenge of real-time learning by integrating Bayesian inference into maintenance optimization to dynamically update beliefs about component degradation parameters. While promising, such Bayesian models have primarily been studied in single-asset contexts [Elwany et al., 2011, Kim and Makis, 2013, Chen et al., 2015, Van Oosterom et al., 2017, Drent et al., 2023], where they remain analytically tractable.
Recent work in the field of maintenance optimization has made progress either in (i) integrating uncertain prognostics into maintenance planning [Mitici et al., 2023, Zhuang et al., 2023] and developing maintenance planning under partial observability or model uncertainty [Arcieri et al., 2023, Tseremoglou and Santos, 2024], or (ii) optimizing economically dependent multi-component CBM under a known degradation model [Zhang et al., 2024]. What remains largely missing is a unified approach in which parameter uncertainty evolves over time and is updated as new degradation observations arrive, while policies still coordinate actions across assets due to shared setup costs—such as the fixed expenses incurred when dispatching maintenance engineers to service equipment.
In this paper, we contribute: (i) a CBM model for economically dependent assets with unknown and heterogeneous degradation parameters; (ii) a dual formulation as both a partially observable Markov decision process (POMDP) and, under conjugacy assumptions, a Bayesian Markov decision process (BMDP) with a compact belief-state representation; (iii) two solution approaches based on deep reinforcement learning (DRL) for the POMDP/BMDP maintenance optimization problem: (a) open-loop deployment of policies trained under full information and (b) direct training on posterior information; (iv) analytical monotonicity results for the full-information Markov decision process (MDP) with respect to both degradation levels and parameters; and (v) numerical and case-study evidence quantifying when belief-aware training is essential and when open-loop deployment degrades under high heterogeneity.
@@
The assets are assumed to degrade independently, with the shared setup cost constituting the only dependence between assets in the network. In addition, costs are incurred for corrective or preventive component replacements. Other factors, such as safety hazards or operational disruptions, are often incorporated indirectly through the corrective-to-preventive cost ratio [De Jonge and Scarf, 2020]. Component degradation is modeled using stochastic shock models [Esary and Marshall, 1973, Sobczyk, 1987]. We assign prior distributions to represent uncertainty about key degradation parameters, specifically, the shock occurrence rate and the distributional parameter for the damage incurred per shock. Conjugate priors facilitate analytically tractable Bayesian updating as degradation data accumulate, following the single-asset framework in [Drent et al., 2023].
@@
To address the resulting high-dimensional control problem, we develop methodological innovations and insights that enable the training of DRL-based policies to optimize CBM for networks of interdependent assets that degrade according to a shock process with unknown parameters, which are progressively inferred from real-time degradation signals. Specifically, we develop two distinct Bayesian simulation environments for a multi-asset setting, each supporting a different DRL approach. The first modeling approach explicitly simulates the true degradation parameters for each component; these are drawn from a population distribution at the time of component replacement. Subsequent degradation then evolves according to a stochastic shock process governed by these sampled parameters. Since the true degradation parameters are unobservable in practice, this setup constitutes a POMDP. To enable deployment of the trained policies, we propose an open-loop feedback approach. Concretely, we collapse the posterior distributions into point estimates, which serve as input to the neural network in place of the true degradation parameters. To complement this first approach, we propose a second modeling approach that assumes the availability of a conjugate prior and, in that case, is equivalent in objective to the first approach—although the two differ in methodology. Instead of simulating the true degradation parameters, this approach explicitly maintains and updates uncertainty by tracking only the posterior distributions of the parameters. These posteriors are continuously updated through real-time degradation observations. This approach constitutes a BMDP, which forms the foundation for the development of DRL-based policies that operate directly on the posterior distributional information available to the asset manager. This marks a significant advancement beyond single-asset models by jointly addressing learning and maintenance optimization in heterogeneous asset networks.
@@
As the training algorithm for our DRL-based policies, we employ specific forms of approximate policy iteration (API) [Temizöz et al., 2025], in which the derived monotonicity results are incorporated both through action space constraints during training and in the design of the initialization heuristics. In particular, two heuristic policies are adopted as benchmarks: (i) a two-threshold control-limit heuristic that groups maintenance actions and is used to initialize API, and (ii) an integrated Bayes heuristic known to yield near-optimal performance in the single-asset setting.
@@
Through extensive numerical experiments and a practical case study utilizing real-world degradation data from interventional X-ray (IXR) filaments, we demonstrate the effectiveness and practical applicability of our DRL approaches. To better understand how policy performance varies with the availability of degradation information, we extend the concept of information levels [Da Costa et al., 2023] to our Bayesian setting: Information level $\textbf{L}_{2}$ corresponds to full knowledge of the degradation parameters for each component. In information level $\textbf{L}_{1}$, this parameter information is unknown, but the distribution of degradation parameters is available. Finally, information level $\textbf{L}_{0}$—used only in the case study—represents a setting where even the distributional information must be estimated from historical data. Although the proposed framework is in this paper directly applied to medical equipment—specifically the filament component of the IXR system—it has broad applicability across other domains, including wind farms, wafer steppers, manufacturing processes, and beyond. The integration of Bayesian parameter learning, economic dependence, and DRL-based policy optimization applies to asset networks where degradation parameters are unknown and need to be learned over time.
@@
The remainder of the paper is organized as follows. Section 2 presents an overview of the relevant literature. We provide a detailed model formulation in Section 3. In Section 4, we detail the heuristic solutions and provide a brief overview of the DRL algorithm. We demonstrate the effectiveness of the proposed solutions through a concise simulation study in Section 5. In Section 6, we present the results of the case study on degrading IXR filaments. We offer concluding remarks in Section 7. A visual roadmap of the paper’s structure and the relationships between its main components is provided in Figure 1.
Figure 1: Flowchart illustrating the conceptual structure of the paper and the relationships between the modeling framework, theoretical results, algorithmic design, and numerical results.
@@ 2 Literature review
Our work focuses on the optimization of CBM for asset networks with economic dependencies, the learning of maintenance strategies under component heterogeneity, and the application of DRL to CBM. We next review these three research streams.
@@
Maintenance models may be classified as single-asset or multi-asset [Olde Keizer et al., 2017, De Jonge and Scarf, 2020]. Multi-asset models extend single-asset models by incorporating joint maintenance policies, accounting for various dependencies such as economic, structural, stochastic, or resource dependencies. Economic dependence refers to the cost relationship between joint maintenance of multiple assets and the maintenance of individual assets. If joint maintenance is more expensive, it indicates negative economic dependence; if it is less expensive, it indicates positive economic dependence.
@@
Various models for CBM in systems with economic dependence have been developed in recent literature due to their relevance in industrial settings. Wijnmalen and Hontelez [1997] propose a heuristic algorithm that minimizes long-run average costs in a multi-component system by determining simple, component-wise control limits and leveraging coordinated repairs to reduce shared setup costs. Bouvard et al. [2011] propose a maintenance optimization method that dynamically groups tasks for multi-component systems using updated degradation data, demonstrated on commercial heavy vehicles. Tian et al. [2011] propose a CBM strategy for wind farms that leverages neural networks to predict the degradation level and corresponding remaining useful lifetime (RUL) from real-time data. Tian and Liao [2011] develop a CBM policy for multi-component systems using a proportional hazards model and a numerical algorithm to evaluate the cost of the proposed policy.
@@
Zhu et al. [2015] introduce a novel CBM policy for multi-component systems experiencing continuous stochastic degradation, utilizing a joint maintenance interval to reduce shared setup costs. Olde Keizer et al. [2016] develop a dynamic programming model to determine the optimal CBM strategy for systems with both economic dependencies and redundancy, demonstrating that it significantly outperforms heuristic policies and offers insights into the optimal policy structure. Olde Keizer et al. [2018] investigate the trade-off between prompt replacement of failed components and maintenance clustering opportunities in redundant systems with economic dependencies, revealing that heuristic threshold policies may be suboptimal and that misinterpreting load-sharing effects can lead to higher maintenance expenses. Do et al. [2019] present a model for a CBM policy in a two-component system with stochastic dependencies, where the degradation of each component is influenced by the other’s state, and economic dependencies, where combined maintenance activities are shown to be more cost-effective. Similarly, Oakley et al. [2022] propose a CBM policy for multi-component systems that balances the trade-offs between the urgency of replacing failed components due to increased load and the economic benefits of clustering replacements to minimize downtime and maintenance costs. Zhang et al. [2024] incorporates imperfect maintenance into CBM optimization for multi-component systems using genetic algorithms. We refer the reader to Zhu et al. [2015, Table 1] for a more extensive summary of CBM models proposed for multi-component systems.
@@
The geographical layout of asset networks often introduces a complex dependency. Abdul-Malak and Kharoufeh [2018] investigate the challenge of optimally replacing multiple stochastically degrading systems within a shared environment using CBM. Soltani et al. [2024] study the problem of optimally maintaining an offshore wind turbine farm under both economic and stochastic dependence due to shared maintenance setup costs and their common environment. They establish monotonicity of the cost function jointly in the degradation level and environmental state, characterize the structure of the optimal replacement policy, and show that sharing maintenance resources is cost-effective. Leppinen et al. [2025] introduce directed graphs to represent the economic and structural dependencies of a multi-component system, including scenarios where maintenance on one component may require the disassembly or maintenance of others, and solve the resulting MDP using a modified policy iteration algorithm to determine the most cost-efficient policy.
@@
Most existing studies on uncertainty in degradation processes typically assume a specific parametric form, with uncertainty modeled through its parameters and captured via a prior distribution. Bayesian inference is then used to update this uncertainty over time—but so far, these approaches have been limited to single-asset systems only. Van Oosterom et al. [2017] develop a POMDP model to incorporate population heterogeneity for maintenance scheduling of a single-asset system that stochastically degrades; however, the population of spare parts consists of multiple indistinguishable types that degrade at varying rates. Elwany et al. [2011] and Si et al. [2018] both study a Wiener degradation process with an unknown drift parameter. In both studies, periodic inspections are used to estimate the parameter, with Bayesian inference applied in the former and maximum likelihood estimation in the latter. These estimation methods, combined with observed degradation levels, guide the decision-making process for maintenance interventions. Flage et al. [2012] consider a single asset subject to a degradation process with unknown parameters. Sequential inspections assess the degradation level, with the timing of the next inspection based on the current degradation level. Uncertainty in the parameters is modeled using a prior distribution and updated in a Bayesian manner. Drent et al. [2023] investigate a single-asset model with one critical component that degrades according to a compound Poisson process with unknown parameters, leveraging real-time degradation data to infer the component’s degradation behavior and adjust decision-making accordingly. They demonstrate that, accounting for component heterogeneity, the optimal policy depends on both the asset’s age and the observed degradation signal, and that integrating learning with decision-making yields significant cost reductions.
In contrast to Drent et al. [2023] and other single-asset models described above, our work considers a network of economically dependent assets, where managers must account for shared setup costs and coordinate maintenance decisions across multiple assets. Extending such single-asset models to this multi-asset setting substantially increases computational complexity, rendering exact solution approaches for the resulting BMDP intractable. We therefore develop scalable DRL-based policies that enable integrated learning and maintenance optimization in heterogeneous asset networks under parameter uncertainty.
@@
DRL-based approaches to maintenance optimization have demonstrated promising results across various settings. Kuhnle et al. [2019] train opportunistic DRL-based maintenance strategies using proximal policy optimization for parallel assets, achieving reductions in downtime and costs compared to traditional strategies such as reactive and time-based maintenance. Zhang and Si [2020] propose a DRL algorithm based on deep Q-network (DQN) to train policies for CBM in multi-component systems with stochastic and economic dependencies, and show in a numerical study that the trained policies outperform benchmarks set by heuristic policies. Mohammadi and He [2022] develop a DRL-based method using double DQN (DDQN) for the joint optimization of maintenance and renewal planning under practical constraints. Using historical inspection and maintenance data to simulate a rail infrastructure environment, they show that the approach generates cost-effective policies that enhance network reliability and safety. Hung et al. [2024] employ DDQN to train DRL-based maintenance policies in a stochastic factory setting characterized by various degradation levels, uncertain repair times and fluctuating machine workloads.
Lee and Mitici [2023] develop an integrated predictive maintenance framework that combines RUL estimation via convolutional neural networks with a DRL-based maintenance policy trained using a soft actor–critic algorithm. The policy triggers maintenance actions based on the estimated RUL distribution and yields significant cost savings and downtime reduction for aircraft turbofan engines. Tseremoglou and Santos [2024] propose a two-stage CBM framework for aircraft fleets, where RUL predictions are first used to construct a maintenance policy via a POMDP and then integrated into a rolling-horizon DQN to schedule preventive and corrective tasks subject to resource constraints. Zhuang et al. [2023] proposes a prognostics-driven framework that uses Bayesian deep learning to generate uncertainty-aware RUL distributions and dynamically updates maintenance and spare-part ordering decisions for turbofan engines.
@@
Da Costa et al. [2023] propose a DRL approach based on distributional DQN to minimize maintenance and downtime costs in asset networks serviced by a single maintenance engineer, where degradation levels are only partially observable. Building upon this work, Verleijsdonk et al. [2024] develop DRL-based solutions trained via a form of API—shown to scale better than the aforementioned training algorithms—for industrial asset networks maintained by multiple engineers; however, their approach assumes full observability of all degradation levels.
@@
Recent studies have made important progress (see, e.g., Andriotis and Papakonstantinou [2021]) by integrating Bayesian inference with DRL in the context of maintenance optimization. These efforts focus on inferring the degradation state of components rather than the parameters of the underlying degradation process. Thus, while such studies represent key steps toward maintenance optimization under uncertainty, they typically assume that the parameters of the degradation processes are known and can be used to define the simulation environment in which the DRL-based policy is trained.